Fly-By-Night Couriers is analyzing the possible acquisition of Flash-in-the-Pan Restaurants. Neither firm has debt. The forecasts of Fly-By-Night show that the purchase would increase its annual aftertax cash flow by $390,000 indefinitely. The current market value of Flash-in-the-Pan is $11 million. The current market value of Fly-By-Night is $27 million. The appropriate discount rate for the incremental cash flows is 8 percent. Fly-By-Night is trying to decide whether it would offer 35 percent of its stock or $14 million in cash to Flash-in-the-Pan.

a.

What is the synergy from the merger? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars,  rounded to the nearest whole number, e.g., 1,234,567.)

b. What is the value of Flash-in-the-Pan to Fly-By-Night? (Do not round intermediate calculations and enter your answer in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567.)
c. What is the cost to Fly-By-Night of each alternative? (Do not round intermediate calculations and enter your answers in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567.)
d. What is the NPV to Fly-By-Night of each alternative? (Do not round intermediate calculations and enter your answers in dollars, not millions of dollars, rounded to the nearest whole number, e.g., 1,234,567.)
e. What alternative should Fly-By-Night use?
Answer #1

(a): Synergy = present value of the incremental cash flows of the proposed purchase

= 390,000/8%

= $4,875,000

(b): Value = synergy + current market value of Flash in the pan

= 4,875,000 + 11,000,000

= $15,875,000

(c ): Cost of cash alternative = 14,000,000 – 11,000,000

= $3,000,000

Cost of stock alternative = 0.35*(15,875,000+27,000,000) – 11,000,000

= $4,006,250

(d): NPV = Value of Flash-in-the-Pan to Fly-by-Night – (equivalent) cash offer = synergy – cost

NPV of cash alternative = 4,875,000 – 3,000,000

= $1,875,000

NPV of stock alternative = 4,875,000 – 4,006,250

= $868,750

(e ): Use cash alternative as its NPV is higher.